# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ Acknowledgements.
L.M.’s research is supported by the ANR grant ProGraM (ANR-19-CE40-0025). V.P. acknowledges the support of the ANR through the grant DOOM (ANR-23-CE23-0002)
@@ 2 Unimodular matched random graphs
The goal of this section is to define precisely the objects we are working on, and the precise notions of optimality for matchings and local convergence we will use. The key concept is unimodularity for rooted graphs. Depending on the context, it is often easier to work either with graphs rooted at a vertex or with graphs rooted at an oriented edge. We will present both cases, as well as the classical correspondence between the two points of view. For a more complete exposition on the subject, we refer the reader to [aldous2018processes].
Finally, we will focus on unimodular matchings and a concept of optimality with respect to a natural performance for vertex-rooted or edge-rooted graphs. We will see that the correspondence mentioned above conserves our notions of optimality.
2.1 Unimodular rooted graphs
@@
Let $G=(V,E)$ be a (locally finite) graph and $w$ be a function from $E$ to $\mathbb{R}$. We that $(G,o,w)$ is
  • a vertex-rooted weighted graph if $o\in V$,
  • an edge-rooted weighted graph if $o\in\overset{\rightarrow}{E}$, the set of directed edges.
It will be useful to add decorations to our graphs: we allow the existence of additional functions $(f_{i})_{i\in\{1..I\}}$ that map $E$ or $\overset{\rightarrow}{E}$ to $\mathbb{R}$. The integer $I$ will typically be no larger than $4$.
@@
Let $(G,o,w)$ be a vertex-rooted (resp. edge-rooted) weighted graph. Fix $I$ an integer and $(f_{i})_{i\in\{1...I\}}$ some functions from $E$ or $\overset{\rightarrow}{E}$ to $\mathbb{R}$. We say that $(G,o,w,(f_{i})_{i\in\{1,...I\}})$ is a decorated vertex-rooted (resp. edge-rooted) weighted graph.
Let us name $\mathcal{G}^{*}$ the space of locally finite decorated vertex-rooted weighted graphs. We will write $\mathcal{L}(\mathcal{G}^{*})$ for the space of laws on this space. Similarly, we denote by $\hat{\mathcal{G}}^{*}$ and $\mathcal{L}(\hat{\mathcal{G}}^{*})$ the corresponding edge-rooted space and laws. When we do not need to keep track of the weights, root, or some decorations of the graphs, we will denote elements of $\mathcal{G}^{*}$ indifferently by $G$, $(G,o)$, $(G,o,w)$, $(G,o,w,(f_{i})_{i\in\{1,...I\}})$, …, keeping only the quantities we are currently interested in.
Remark.
The labels of the vertices being of no importance we work up graph isomorphism.
We now introduce the notion of unimodularity, which translates the intuition that the choice of the root is uniform for some random graph $\mathbb{G}$ even when the graph is infinite. To this end, we introduce the space of doubly rooted graphs up to isomorphism $\mathcal{G}^{**}$ which is defined similarly as before but with two distinguished roots. It will be simpler to give specific definitions for vertex and edge-rooted graphs.
@@
We say that a probability measure $\mu$ on decorated vertex-rooted graphs is unimodular if the following statement holds for every $f:\mathcal{G}^{**}\mapsto\mathbb{R}_{+}$:
$\int_{\mathcal{G}^{*}}\left(\sum_{v\in V(G)}f(G,o,v)\right)\,\mathrm{d}\mu(G,o)=\int_{\mathcal{G}^{*}}\left(\sum_{v\in V(G)}f(G,v,o)\right)\,\mathrm{d}\mu(G,o).$
The subspace of unimodular laws on vertex rooted-graph will be noted $\mathcal{L}_{U}(\mathcal{G}^{*})$
@@
The subspace of unimodular laws on edge-rooted graphs will be noted $\mathcal{L}_{U}(\hat{\mathcal{G}}^{*})$
@@
Fix $G$ a finite deterministic graph, and let $\mathbf{o}$ be a random vertex (resp. oriented edge). It is straightforward to check that $(G,\mathbf{o})$ is unimodular iff $\mathbf{o}$ is uniform. Hence, unimodular graphs can be viewed as generalisations of uniformly rooted graphs.
2.2 Correspondance between vertex-rooted and edge-rooted unimodularity
In this section, we present the classical transformation that maps a unimodular vertex-rooted graph into a unimodular edge-rooted graph. Heuristically, to transform a graph rooted at a uniform vertex into a graph rooted at a uniformly oriented edge, one has to pick an oriented edge starting at its root vertex, but this induces a bias by the degree of the root vertex. This bias has to be taken into account to conserve unimodularity. This is done in the following.
@@ 2.2 Correspondance between vertex-rooted and edge-rooted unimodularity
Take $\mu_{v}\in\mathcal{L}(\mathcal{G}^{*})$ such that $m=\int_{\mathcal{G}^{*}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$. Let $R(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{*})$ be the unique measure such that, for every $f$ measurable from $\hat{\mathcal{G}}^{*}$ to $\mathbb{R}_{+}$,
$\int_{\hat{\mathcal{G}}^{*}}f(G,(o_{-},o_{+}))\,\mathrm{d}R(\mu_{v})(G,(o_{-},o_{+}))=\int_{\mathcal{G}*}\frac{1}{m}\sum_{u\sim o}f(G,o,u)\,\mathrm{d}\mu_{v}(G,o).$
@@
The operator $R$ is the composition of two transformations. The first operator consists on choosing a uniform oriented edge started at the root vertex, giving a measure $R_{1}(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{*})$:
$\int_{\hat{\mathcal{G}}^{*}}f(G,(o_{-},o_{+}))\,\mathrm{d}R_{1}(\mu_{v})(G,(o_{-},o_{+})))=\int_{\mathcal{G}*}\frac{1}{\deg(o)}\sum_{u\sim o}f(\textbf{}G,o,u)\,\mathrm{d}\mu_{v}(G,o).$
The second step is then to cancel the biais by the degree of the vertex, giving $R(\mu_{v})=R_{2}\circ R_{1}(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{*})$:
$\int_{\hat{\mathcal{G}}^{*}}f(G,(o_{-},o_{+}))\mathrm{d}R_{2}\circ R_{1}(\mu_{v})(G,(o_{-},o_{+}))=\int_{\hat{\mathcal{G}}^{*}}\frac{\deg(o_{-})}{m}f(G,(o_{-},o_{+}))\,\mathrm{d}R_{1}(\mu_{v})(G,(o_{-},o_{+})).$
@@
Proposition 5 (Theorem 4.1 in [aldous2018processes]).
Let $\mu_{v}\in\mathcal{L}(\mathcal{G}^{*})$ such that $m=\int_{\mathcal{G}^{*}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$. Then the measure $\mu_{v}$ is unimodular if and only if $R(\mu_{v})$ is unimodular.
We finish this Section with the classical example of unimodular Bienaymé-Galton-Watson trees (UGBW), seen from both points of view and without weights. In either case, the weighted version with weight law $\omega$ corresponds to drawing $(w(e))_{e\in E}$ independently of law $\omega$.
Example 2.1 (Unimodular Bienaymé Galton Watson trees)
Let $\pi$ be a probability measure on =$\mathbb{N}=\{0,1,...\}$ with finite expectation $m>0$. Let $\hat{\pi}$ be the size-biased version of $\pi$, that, $\forall k\geq 0$, $\hat{\pi}(k)=\frac{k}{m}\pi(k)$.
In the introduction, we defined the Unimodular Bienaymé Galton Watson Tree (what we will shorten as UBGW tree) with reproduction law $\pi$. It is the random tree $\mathbb{T}$ with the following law:
@@
The corresponding edge-rooted UBGW tree has a very simple description. Take two independent copies of Bienaymé-Galton-Watson trees with offspring distribution $\hat{\pi}$ with respective root vertex $o_{-}$ and $o_{+}$, and connect their roots by the oriented edge $(o_{-},o_{+})$. The resulting random tree $(\hat{\mathbb{T}},(o_{-},o_{+}))$ is an edge-rooted unimodular random graph. See Figure 2 for an illustration.
$o_{-}$$o_{+}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$
Note that UGBW trees are the limits in law for the local topology of random configuration graphs with iid degrees sampled for $\pi$, and rooted uniformly at a vertex or a directed edge.
@@ 2.3 Matchings, optimality and unimodularity
For any weighted graph $G=((V,E),w)$, a matching $M=(V^{\prime},E^{\prime})$ on $G$ is a subgraph of $G$ in the sense $V=V^{\prime}$ and $E^{\prime}\subset E$, such that every vertex of $V$ belongs to at most one edge of $E^{\prime}$. A matched graph is a pair $(G,M)$, where $M$ is a matching on $G$.
We can extend the notion of unimodularity for random graphs to random matched graphs by assimilating a matching to a decoration.
@@
Let $(\mathbb{G},\mathbb{M},\mathbf{o})$ be a random matched rooted graph. We say that $(\mathbb{G},\mathbb{M},\mathbf{o})$ is unimodular iff $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})$ is. By convention, we will append the matching as the final decoration.
Remark.
Note that this definition does not require that the matching is measurable with respect to the graph. This will be important in the sequel.
The central objects of this work are optimal matchings. For finite graphs, this corresponds to the classical notion of maximum weighted matching. When $G$ is infinite, this optimality is ill-defined. However, in the case of a unimodular random weighted graph, since the root edge is informally a typical edge, we can define optimality via its weight when it is in the matching. This leads to the following definition:
@@
Let $(\mathbb{G},\mathbf{o},\mathbb{M},)$ be a unimodular random matched rooted graph, we define the performance of $(\mathbb{G},\mathbf{o},\mathbb{M})$ as:
@@
When the context is clear, we will just shorten the notation to $\operatorname{perf}(\mathbb{M})$. By extension, since those quantities only depend on the law of $(\mathbb{G},\mathbf{o},\mathbb{M})$, we will freely use the same notation $\operatorname{perf_{E}}(\mu_{o,m}):=\operatorname{perf_{E}}(\mathbb{G},\mathbb{M})$ for $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})\sim\mu_{o,m}$ and $\operatorname{perf_{V}}(\mu_{v,m}):=\operatorname{perf_{V}}(\mathbb{G},\mathbb{M})$ for $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})\sim\mu_{v,m}$ where $\mu_{e,m}\in\mathcal{L}_{U}(\hat{\mathcal{G}}^{*})$ satisfies that the last decoration defines a matching almost surely and the same for $\mu_{v,m}\in\mathcal{L}_{U}(\mathcal{G}^{*})$.
Let $(\mathbb{G},\mathbf{o})$ be a (undecorated) unimodular vertex-rooted graph. We say that $(\mathbb{G}^{\prime},\mathbf{o},\mathbb{M})$ is optimal if it is a unimodular matched vertex-rooted graph such that
  • $(\mathbb{G}^{\prime},\mathbf{o}^{\prime})$ has the same law as $(\mathbb{G},\mathbf{o})$
  • $\operatorname{perf_{V}}((\mathbb{G}^{\prime},\mathbf{o}^{\prime}),\mathbb{M})$ is maximal.
@@
This formulation may seem a bit complicated but it is possible that $\mathbb{M}$ is not measurable with respect to $(\mathbb{G},\mathbf{o})$. We thus allow resampling to avoid dealing with this difficulty.
In the edge-rooted setting, we can define optimality similarly. The next proposition shows that the operator $R$ preserves optimality.
Proposition 6.
Let $\mu_{v,m}\in\mathcal{L}_{U}(\mathcal{G}^{*})$ such that the last decoration defines a matching almost surely, assume $m=\int_{\mathcal{G}^{*}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$. Recall the operator $R$ of Definition 2.5, then:
$\operatorname{perf_{V}}(\mu_{v,m})=m\operatorname{perf_{E}}(R(\mu_{v,m})).$
@@
If the graph is finite and the root is chosen uniformly among the vertices or the directed edges, then $\operatorname{perf_{V}}$ is simply the average contribution per vertex, and $\operatorname{perf_{E}}$ is the average contribution per directed edge. It is then clear that the two quantities are proportional. The proposition shows that it generalises to unimodular matched graphs.
@@
We will decompose both quantities with respect to the degree of the root, let us write $(\mathbb{G},o,\mathbbm{1}_{\mathbb{M}})\sim\mu_{v,m}$ and $(\hat{\mathbb{G}},(o_{-},o_{+}),\mathbbm{1}_{\hat{\mathbb{M}}})\sim R(\mu_{v,m})$. In the edge-rooted case, we get:

### v2

@@ Acknowledgements.
L.M.’s research is supported by the ANR grant ProGraM (ANR-19-CE40-0025). V.P. acknowledges the support of the ANR through the grant DOOM (ANR-23-CE23-0002) and through the PEPR IA FOUNDRY project (ANR-23-PEIA-0003).
@@ 2 Unimodular matched random graphs
The goal of this section is to precisely define our objects of interest. The key concepts are local topology and unimodularity of rooted graphs. Depending on the context, it is often easier to work with graphs rooted either at a vertex or at an oriented edge. We will present both cases, as well as the classical correspondence between the two points of view. For a more complete exposition on the subject, we refer to [aldous2018processes].
Finally, we will define unimodular matchings and a concept of optimality with respect to a natural performance for vertex-rooted or edge-rooted graphs. We will see that the correspondence mentioned above conserves our notion of optimality.
2.1 Rooted graphs and local convergence
@@
Let $G=(V,E)$ be a (locally finite) graph. We call vertex-rooted weighted graph (resp. edge-rooted weighted graph), the triplet $(G,o,w)$, where $o\in V$ (resp. $o\in\overset{\rightarrow}{E}$) and $w$ is a function from $E$ to $\mathbb{R}$.
It will be useful to add graph decorations, namely functions that map $\overset{\rightarrow}{E}$ to $\overline{\mathbb{R}}$.
@@
Let $(G,o,w)$ be a vertex-rooted (resp. edge-rooted) weighted graph. Fix $I$ an integer and $(f_{i})_{i\in\{1,...,I\}}$ some functions from $\overset{\rightarrow}{E}$ to $\overline{\mathbb{R}}$. We then say that $(G,o,w,(f_{i})_{i\in\{1,...,I\}})$ is a decorated vertex-rooted (resp. edge-rooted) weighted graph. If $f_{i}$ is symmetrical then we will identify it with a map from $E$ to $\overline{\mathbb{R}}$.
Since we are not interested in the labels of the vertices, we will work up to graph isomorphism. We will say that $(G,o,w,(f_{i})_{i\in\{1,...,I\}})\simeq(G^{\prime},o^{\prime},w^{\prime},(f_{i}^{\prime})_{i\in\{1,...,I\}})$ if there exists some one-to-one function $g$ called graph isomorphism from $V$ to $V^{\prime}$ such that if $g(o)=o^{\prime}$ and for all $(u,v)\in V$, $w^{\prime}(g(u),g(v))=w(u,v)$ and for all $i\in\{1,...,I\}$, $f^{\prime}_{i}(g(u),g(v))=f_{i}(u,v)$
Definition 2.3.
Let $\mathcal{G}^{\star}$ be the space of locally finite decorated vertex-rooted weighted graphs up to isomorphism. We will write $\mathcal{L}(\mathcal{G}^{\star})$ for the space of laws on this space. Similarly, we denote by $\hat{\mathcal{G}}^{\star}$ and $\mathcal{L}(\hat{\mathcal{G}}^{\star})$ the corresponding edge-rooted space and laws.
When we do not need to keep track of the weights, root, or some decorations of the graphs, we will denote elements of $\mathcal{G}^{\star}$ indifferently by $G$, $(G,o)$, $(G,o,w)$, $(G,o,w,(f_{i})_{i\in\{1,...,I\}})$, …, keeping only the quantities we are currently interested in.
The topology of local convergence on $\mathcal{G}^{\star}$ and $\hat{\mathcal{G}^{\star}}$ was first introduced by Benjamini-Schramm [BenjaminiSchramm1] and by Aldous-Steele [Aldous2004], whose precise definitions follow. We denote by $d_{gr}$ the graph distance between vertices of a graph. For any vertex-rooted decorated weighted graph $(G=(V,E),o,w,(f_{i})_{i\in I})$ and $H>0$, the $H$-neighborhood $B_{H}((G,o,w,(f_{i})_{i\in\{1,...,I\}}))$ of $o$ in the graph is the vertex-rooted decorated weighted graph $(G_{H},o,w_{H},(f_{i,H})_{i\in\{1,...,I\}})$ with $G_{H}:=(V_{H},E_{H})$ such that
$V_{H}$ =$=\{v\in V,d_{gr}(o,v)\leq H\},$
$E_{H}$ =$=\{(u,v)\in E,(u,v)\in V_{H}^{2}\},$
$w_{H}$ are the weights of edges in $E_{H}$ and, for all $i\in\{1,...,I\}$, the decoration $f_{i,H}$ is the decoration $f_{i}$ restricted to $\overset{\rightarrow}{E}_{H}$.
Let $(G,o,w,(f_{i})_{i\in\{1,...,I\}})$ and $(G^{\prime},o^{\prime},w^{\prime},(f_{i}^{\prime})_{i\in\{1,...,I\}}$ be two vertex-rooted decorated weighted graphs. Let $H\geq 0$ be the largest integer such that there exists a graph isomorphism $g$ from $B_{H}(G,o)$ to $B_{H}(G^{\prime},o^{\prime})$ such that $\|w^{\prime}_{H}\circ g-w_{H}\|_{\infty}\leq\frac{1}{H}$ and $\|f^{\prime}_{i,H}\circ g-f_{i,H}\|_{\infty}\leq\frac{1}{H}$ for all $i\in\{1,\ldots,I\}$. We set
$d_{\mathrm{loc}}\left((G,o,w,(f_{i})_{i\in\{1,...,I\}}),(G^{\prime},o^{\prime},w^{\prime},(f_{i}^{\prime})_{i\in\{1,...,I\}}\right):=\frac{1}{1+H}.$ (8)
The function $d_{\mathrm{loc}}$ is a distance on $\mathcal{G}^{\star}$ and the space $(\mathcal{G}^{\star},d_{\mathrm{loc}})$ is a Polish space. The topology induced by $d_{\mathrm{loc}}$ is called the local topology on $\mathcal{G}^{\star}$. Weak convergence on $\mathcal{L}(\mathcal{G}^{\star})$ for the local topology is called local weak convergence. Similar definitions can be given for the edge-rooted versions.
2.2 Vertex-rooted and edge-rooted unimodularity
We now introduce the notion of unimodularity. To this end, we define the space of doubly rooted graphs up to isomorphism $\mathcal{G}^{\star\star}$ similarly as before, but with two distinguished roots. It will be simpler to give separately specific definitions for vertex-rooted and edge-rooted graphs. We refer to Aldous-Lyons [aldous2018processes] for a comprehensive exposition on the topic.
@@ 2.2 Vertex-rooted and edge-rooted unimodularity
We say that a probability measure $\mu$ on decorated vertex-rooted graphs is unimodular if the following statement holds for every measurable $f:\mathcal{G}^{\star\star}\mapsto\mathbb{R}_{+}$:
$\int_{\mathcal{G}^{\star}}\left(\sum_{v\in V(G)}f(G,o,v)\right)\,\mathrm{d}\mu(G,o)=\int_{\mathcal{G}^{\star}}\left(\sum_{v\in V(G)}f(G,v,o)\right)\,\mathrm{d}\mu(G,o).$
The subspace of unimodular laws on vertex rooted-graph will be noted $\mathcal{L}_{U}(\mathcal{G}^{\star})$
@@
The subspace of unimodular laws on edge-rooted graphs will be noted $\mathcal{L}_{U}(\hat{\mathcal{G}}^{\star})$
@@
Fix $G$ a finite deterministic graph, and let $\mathbf{o}$ be a random vertex (resp. oriented edge). It is straightforward to check that $(G,\mathbf{o})$ is unimodular iff $\mathbf{o}$ is uniform. Hence, unimodular graphs can be viewed as generalisations of uniformly rooted graphs, which is one of the fundamental findings of [BenjaminiSchramm1].
We now present the classical transformation that maps a unimodular vertex-rooted graph to a unimodular edge-rooted graph. Heuristically, to transform a graph rooted at a uniform vertex into a graph rooted at a uniformly oriented edge, one has to pick an oriented edge starting at its root vertex, but this induces a bias by the degree of the root vertex. Indeed, a directed edge $(u,v)$ will be less likely to be the new root the greater the degree of $u$ is. This bias has to be taken into account to conserve unimodularity. This is done in the following transformations on the laws.
@@
Take $\mu_{v}\in\mathcal{L}(\mathcal{G}^{\star})$ such that $0<m:=\int_{\mathcal{G}^{\star}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$. Let $R(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{\star})$ be the unique measure such that, for every $f$ measurable from $\hat{\mathcal{G}}^{\star}$ to $\mathbb{R}_{+}$,
$\int_{\hat{\mathcal{G}}^{\star}}f(G,(o_{-},o_{+}))\,\mathrm{d}R(\mu_{v})(G,(o_{-},o_{+}))=\int_{\mathcal{G}*}\frac{1}{m}\sum_{u\sim o}f(G,o,u)\,\mathrm{d}\mu_{v}(G,o).$
@@
The operator $R$ is the composition of two transformations. The first operator consists of choosing a uniformly oriented edge started at the root vertex, giving a measure $R_{1}(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{\star})$:
$\int_{\hat{\mathcal{G}}^{\star}}f(G,(o_{-},o_{+}))\,\mathrm{d}R_{1}(\mu_{v})(G,(o_{-},o_{+})))=\int_{\mathcal{G}*}\frac{1}{\deg(o)}\sum_{u\sim o}f(\textbf{}G,o,u)\,\mathrm{d}\mu_{v}(G,o).$
The second step is then to cancel the bias by the degree of the vertex, giving $R(\mu_{v})=R_{2}\circ R_{1}(\mu_{v})\in\mathcal{L}(\hat{\mathcal{G}}^{\star})$:
$\int_{\hat{\mathcal{G}}^{\star}}f(G,(o_{-},o_{+}))\mathrm{d}R_{2}\circ R_{1}(\mu_{v})(G,(o_{-},o_{+}))=\int_{\hat{\mathcal{G}}^{\star}}\frac{\deg(o_{-})}{m}f(G,(o_{-},o_{+}))\,\mathrm{d}R_{1}(\mu_{v})(G,(o_{-},o_{+})).$
Equivalently, to obtain an edge-rooted version of a vertex rooted random graph, one can consider the size-biased version of the original measure $\mathrm{d}\mu_{v}(G,o)$ (namely sample a graph with measure $\frac{\deg(o)}{m}\mathrm{d}\mu_{v}(G,o)$) and then from this sampled graph, starting from the root, select an edge uniformly at random to get the edge measure $dR(\mu_{v})$.
@@
Proposition 2.7 (Theorem 4.1 in [aldous2018processes]).
Let $\mu_{v}\in\mathcal{L}(\mathcal{G}^{\star})$ be such that $0<\int_{\mathcal{G}^{\star}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$. Then the measure $\mu_{v}$ is unimodular if and only if $R(\mu_{v})$ is unimodular.
2.3 Unimodular Bienaymé-Galton-Watson trees
In this subsection, we introduce Unimodular Bienaymé-Galton-Watson trees (UBGW) along with models of random graphs that converge locally in law to these trees. We will present both the vertex and the edge-rooted point of view, without weights. In either case, the weighted version with weight law $\omega$ corresponds to drawing $(w(e))_{e\in E}$ independently of law $\omega$.
Let $\pi$ be a probability measure on +=$\mathbb{Z}_{+}=\{0,1,...\}$ with finite expectation $m>0$. Let $\hat{\pi}$ be the size-biased version of $\pi$, that is, $\forall k\geq 0$, $\hat{\pi}(k)=\frac{k}{m}\pi(k)$.
In the introduction, we defined the edge-rooted Unimodular Bienaymé Galton Watson Tree with reproduction law $\pi$. Let us recall the construction: take two independent copies of Bienaymé-Galton-Watson trees with offspring distribution $\hat{\pi}$ with respective root vertex $o_{-}$ and $o_{+}$, and connect their roots by the oriented edge $(o_{-},o_{+})$. The resulting random tree $(\hat{\mathbb{T}},(o_{-},o_{+}))$ is an edge-rooted unimodular random graph. See Figure 1 for an illustration.
$o_{-}$$o_{+}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$$\hat{\pi}$
The vertex-rooted UBGW tree is the random tree $\mathbb{T}$ with the following law:
@@ 2.3 Unimodular Bienaymé-Galton-Watson trees
The most classical examples of random graphs converging to UBGW trees we consider are sparse Erdős–Rényi and configuration models:
  • Sparse Erdős–Rényi: Introduced in the celebrated paper of Erdős and Rényi [ErdosRenyi], for $c>0$ and $N\geq 1$, the random graph $\mathcal{G}(N,\frac{c}{N})$ is defined on the vertex set $\{1,\ldots,N\}$ with independent edges between vertices with probability $\frac{c}{N}$. Once uniformly rooted, these graphs converge locally when $N$ goes to $\infty$, to a UBGW tree with reproduction law Poisson with parameter $c$.
  • Configuration model: This model was introduced by Bollobás in 1980 [B1980] and can be defined as follows. Let $N\geq 1$ be an integer and let $d_{1},\ldots,d_{N}\in\mathbb{Z}_{+}$ be such that $d_{1}+\cdots+d_{N}$ is even. We interpret $d_{i}$ as a number of half-edges attached to vertex $i$. Then, the configuration model associated to the sequence $(d_{i})_{1\leq i\leq N}$ is the random multigraph with vertex set $\{1,\ldots,N\}$ obtained by a uniform matching of these half-edges. If $d_{1}+\cdots+d_{N}$ is odd, we change $d_{N}$ into $d_{N}+1$ and do the same construction. Now, let $\mathbf{d}^{(N)}$ be a sequence of random variables defined on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$ such that for every $N\geq 1$, $\mathbf{d}^{(N)}=(d_{1}^{(N)},\ldots,d_{N}^{(N)})\in\mathbb{Z}_{+}^{N}$. Furthermore, suppose that there exists $\pi$ a probability measure on $\mathbb{Z}_{+}$ with finite first moment such that
    $\forall k\geq 0,\quad\frac{1}{N}\sum\limits_{j=1}^{N}\mathbbm{1}_{d_{j}^{(N)}=k}\underset{N\rightarrow+\infty}{\longrightarrow}\pi(\{k\}).$
    The sequence of random configuration graphs associated to the $\mathbf{d}^{(N)}$ has asymptotically a positive probability to be simple. In addition, this sequence of random graphs, when uniformly rooted, converges locally in law to the UBGW random tree with offspring distribution $\pi$, see [vanderHofstadBook2] for more details.
@@ 2.4 Matchings, optimality and unimodularity
For any weighted graph $G=((V,E),w)$, a matching $M=(V,E^{\prime})$ on $G$ is a subgraph of $G$ such that $E^{\prime}\subset E$ and every vertex of $V$ belongs to at most one edge of $E^{\prime}$. A matched graph is a pair $(G,M)$, where $M$ is a matching on $G$.
We can extend the notion of unimodularity for random graphs to random matched graphs by transforming a matching into a canonical decoration. If $(G,M)=((V,E),(V,E^{\prime}))$ is a matched graph then we can define the associated decoration $\mathbbm{1}_{M}:E\to\{0,1\}$ with $\mathbbm{1}_{M}(u,v)=1$ if and only if $(u,v)\in E^{\prime}$.
@@
Let $(\mathbb{G},\mathbb{M},\mathbf{o})$ be a random matched rooted graph. We say that $(\mathbb{G},\mathbb{M},\mathbf{o})$ is unimodular iff $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})$ is unimodular.
The correspondence between $(\mathbb{G},\mathbb{M},\mathbf{o})$ and $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})$ gives a representation of the space of unimodular laws on decorated matched graphs as a subspace of unimodular laws on decorated graphs $\mathcal{L}_{U,\mathcal{M}}(\mathcal{G}^{\star})\subseteq\mathcal{L}_{U}(\mathcal{G}^{\star})$.
@@
The subspace $\mathcal{L}_{U,\mathcal{M}}(\mathcal{G}^{\star})\subseteq\mathcal{L}_{U}(\mathcal{G}^{\star})$ is defined as the set of elements $\mu$ of $\mathcal{L}_{U}(\mathcal{G}^{\star})$ such that for $(G,o,w,(f_{i})_{i\in\{1,..,I\}})\sim\mu$, we have that $f_{I}(u,v)$ maps $E$ to $\{0,1\}$ and the subset $\{(u,v),f_{I}(u,v)=1\}$ almost surely induces a matching on $G$ in the sense that
$\mathbb{P}\left(\exists u,u^{\prime},v\in V,u\neq u^{\prime}\text{ and }f(v,u)=f(v,u^{\prime})=1\right)=0.$
Similarly we can define the edge rooted version $\mathcal{L}_{U,\mathcal{M}}(\hat{\mathcal{G}}^{\star})$.
We recall that the central object of interest is optimal matching, which corresponds to the classical notion of maximal weight matching in finite graphs. When $G$ is infinite, this optimality is ill-defined. However, in the case of a unimodular random weighted graph, since the root edge is informally a typical edge, we can define optimality via its expected weight when it belongs to the matching. This leads to the following definition:
Definition 2.11.
Let $(\mathbb{G},\mathbf{o},\mathbb{M})$ be a unimodular random matched rooted graph, we define the performance of $(\mathbb{G},\mathbf{o},\mathbb{M})$ as:
@@
When the context is clear, we will shorten the notation to $\operatorname{perf}(\mathbb{M})$. By extension, since those quantities only depend on the law of $(\mathbb{G},\mathbf{o},\mathbb{M})$, we will freely use the same notation $\operatorname{perf_{E}}(\mu_{e}):=\operatorname{perf_{E}}(\mathbb{G},\mathbf{o},\mathbb{M})$ for $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})\sim\mu_{e}$ and $\operatorname{perf_{V}}(\mu_{v}):=\operatorname{perf_{V}}(\mathbb{G},\mathbf{o},\mathbb{M})$ for $(\mathbb{G},\mathbf{o},\mathbbm{1}_{\mathbb{M}})\sim\mu_{v}$ where $\mu_{e}\in\mathcal{L}_{U,\mathcal{M}}(\hat{\mathcal{G}}^{\star})$ and $\mu_{v}\in\mathcal{L}_{U,\mathcal{M}}(\mathcal{G}^{\star})$.
Let $(\mathbb{G},\mathbf{o})$ be a (undecorated) unimodular vertex-rooted graph. We say that $(\mathbb{G}^{\prime},\mathbf{o}^{\prime},\mathbb{M})$ is optimal if it is a unimodular matched vertex-rooted graph such that $\operatorname{perf_{V}}((\mathbb{G}^{\prime},\mathbf{o}^{\prime}),\mathbb{M})$ is maximal among all unimodular rooted matched random graphs $(\mathbb{G}^{\prime},\mathbf{o}^{\prime},\mathbb{M}^{\prime})$ such that $(\mathbb{G}^{\prime},\mathbf{o}^{\prime})$ has the same law as $(\mathbb{G},\mathbf{o})$.
@@
One could be tempted to look at performance only for objects of the form $(\mathbb{G},\mathbf{o},F(\mathbb{G},o))$ where $F$ deterministically maps vertex-rooted weighted graphs $(V,E,o,w)$ to decorations from $E$ to $\{0,1\}$ that induces a matching. This paper does not investigate whether this restriction is relevant.
In the edge-rooted setting, optimality is defined similarly. The next proposition shows that the operator $R$, introduced in Definition 2.6, preserves optimality.
Proposition 2.12.
Let $\mu_{v}\in\mathcal{L}_{U,\mathcal{M}}(\mathcal{G}^{\star})$, assume $m=\int_{\mathcal{G}^{\star}}\deg(o)\,\mathrm{d}\mu_{v}(G,o)<\infty$, then:
$\operatorname{perf_{V}}(\mu_{v})=m\operatorname{perf_{E}}(R(\mu_{v})).$
@@
If the graph is finite and the root is chosen uniformly either among the vertices or among the directed edges, then $\operatorname{perf_{V}}$ is simply the average contribution per vertex, and $\operatorname{perf_{E}}$ is the average contribution per directed edge. It is then clear that the two quantities are proportional. The proposition shows that it generalises to unimodular matched graphs.
@@
We will decompose both quantities with respect to the degree of the root. Let us write $(\mathbb{G},o,\mathbbm{1}_{\mathbb{M}})\sim\mu_{v}$ and $(\hat{\mathbb{G}},(o_{-},o_{+}),\mathbbm{1}_{\hat{\mathbb{M}}})\sim R(\mu_{v})$. In the edge-rooted case, we get: